Calculate the total resistance of two, three, four and five 60 Ω resistors in parallel. What is the simple relationship for the total resistance of equal resistances in parallel?


1
Expert's answer
2021-12-28T09:38:14-0500

Total resistance of two 60 Ω resistors in parallel

"\frac{1}{R_{T}}=\frac{1}{R_{1}}+\frac{1}{R_{2}}"

"=\frac{1}{60}+\frac{1}{60}"

"=\frac{2}{60}=\frac{1}{30}"

"\frac{1}{R_{T}}=\frac{1}{30}"

"{R_{T}}=30\u03a9"



Total resistance of three 60Ω resistors in parallel

"\frac{1}{R_{T}}=\frac{1}{R_{1}}+\frac{1}{R_{2}}+\frac{1}{R_{3}}"

"=\frac{1}{60}+\frac{1}{60}+\frac{1}{60}"

"=\frac{3}{60}"

"=\frac{1}{20}"

"\frac{1}{R_{T}}=\frac{1}{20}"

"{R_{T}}=20\u03a9" 



Total resistance of four 60Ω resistors in parallel

"\frac{1}{R_{T}}=\frac{1}{R_{1}}+\frac{1}{R_{2}}+\frac{1}{R_{3}}+\frac{1}{R_{4}}"

"=\frac{1}{60}+\frac{1}{60}+\frac{1}{60}+\frac{1}{60}"

"=\frac{4}{60}"

"\frac{1}{R_{T}}=\frac{1}{15}"

"R_{T}=15\u03a9"




Total resistance of five 60Ω resistors in parallel

"\frac{1}{R_{T}}=\frac{1}{R_{1}}+\frac{1}{R_{2}}+\frac{1}{R_{3}}+\frac{1}{R_{4}}+\frac{1}{R_{5}}"

"=\frac{1}{60}+\frac{1}{60}+\frac{1}{60}+\frac{1}{60}+\frac{1}{60}"

"=\frac{5}{60}"

"\frac{1}{R_{T}}=\frac{1}{12}"

"R_{T}=12\u03a9"



The simple relationship for the total resistance of equal resistances in parallel if given a 60Ω resistor is

"R_{T}=\frac{60Ω}{n}" for "n=2,3,4,..."